OpenStudy (paulalovee):

Pleassee helpp <3 The volume of a cylinder, found using 3.14 to approximate pi, is 50,240 cm3. The radius of the cylinder is 40 cm. What is the height of the cylinder? Express your answer as a whole number. cm

OpenStudy (paulalovee):

Or check my answer

jabez177 (jabez177):

What's your answer? @KendrickLamar2014

OpenStudy (paulalovee):

Nope nvm I dont have one ..

jabez177 (jabez177):

Okay, I somewhat forgot how to do this lol... :/ I'll try and solve it.

OpenStudy (paulalovee):

Ty ty <3

OpenStudy (kendricklamar2014):

@jabez177 Did you mean to tag me or @paulalovee ?

jabez177 (jabez177):

You because I forgot how to solve these. xD The previous one I knew but this one I forgot... I'm watching a video: https://www.youtube.com/watch?v=MxaDQNm4cPw

OpenStudy (kendricklamar2014):

\[V = \pi r^2 h\] \[50,240 = (3.14)(40^2)h\]

OpenStudy (kendricklamar2014):

What is 40^2? @paulalovee

OpenStudy (paulalovee):

Thank you guys for your help, @jabez177 tag me in ur question :)

jabez177 (jabez177):

np lol